### v1

# Can Community Notes Replace Professional Fact-Checkers?

Abstract
Two commonly-employed strategies to combat the rise of misinformation on social media are (i) fact-checking by professional organisations and (ii) community moderation by platform users. Policy changes by Twitter/X and, more recently, Meta, signal a shift away from partnerships with fact-checking organisations and towards an increased reliance on crowdsourced community notes. However, the extent and nature of dependencies between fact-checking and helpful community notes remain unclear. To address these questions, we use language models to annotate a large corpus of Twitter/X community notes with attributes such as topic, cited sources, and whether they refute claims tied to broader misinformation narratives. Our analysis reveals that community notes cite fact-checking sources up to five times more than previously reported. Fact-checking is especially crucial for notes on posts linked to broader narratives, which are twice as likely to reference fact-checking sources compared to other sources. In conclusion, our results show that successful community moderation heavily relies on professional fact-checking.
[…]
Can Community Notes Replace Professional Fact-Checkers?
Nadav Borenstein  Greta Warren  Desmond Elliott  Isabelle Augenstein University of Copenhagen nb@di.ku.dkgrwa@di.ku.dk de@di.ku.dkaugenstein@di.ku.dk
1 Introduction
The proliferation of misinformation on social media (Arnold, 2020; Diakopoulos, 2020), along with the rise of generative AI (Augenstein et al., 2024) have led to increasing concerns about its current and future potential harms, (e.g., to health (Clemente et al., 2022)) and threats to democracy and political stability (Reglitz, 2022).
Fact-checkers play a crucial role in combatting misinformation (Graves, 2017), and in recent years, have partnered with social media platforms, e.g., Meta, YouTube, and TikTok, to tackle its spread on these platforms. However, due to the scale of misleading content shared online, community moderation (e.g., options to flag potential misinformation, group/server moderators) is often employed in parallel (Morrow et al., 2022), as a complementary approach (e.g., (Google, 2025); see also the practice of snoping (Pilarski et al., 2024)).
Figure 1: An example of a community note. Notice the fact-checking link and rating.
The expansion of fact-checking projects in the last decade (Lauer and Graves, 2024), alongside their broader initiatives to curb misinformation (e.g., citizen media literacy programmes (Juneja and Mitra, 2022)) have been aided by partnerships with social media platforms such as Meta and Google (Graves and Anderson, 2020), which fund independent fact-checking agencies to fact-check potentially false claims on their platform. [^1] Fact-checkers provide a judgment of claim veracity and exert no influence on the platforms’ content moderation policies (Catalanello and Sanders, 2025).
However, political pressure and accusations of bias and censorship, and most recently, Meta’s announcement of its plans to end its partnerships with fact-checkers in the U.S. and implement a community moderation model (Meta, 2025), threatens the financial stability of fact-checking organisations, and hence, their ability to keep up with the increasing volume and sophistication of misinformation spread (Stencel et al., 2024; IFCN, 2024).
Meta’s recent policy shift also implies that these two strategies (fact-checking and community notes) are independent and in opposition, rather than two complementary strategies of tackling online misinformation. In this paper, we examine Twitter/X community notes as a case study to understand how fact-checking is used in community notes. Specifically, we investigate the following two questions: (RQ1) To what extent do community notes rely on the work of professional fact-checkers? and (RQ2) What are the traits of posts and notes that rely on fact-checking sources? Studying the relationship between fact-checking and community notes is vital for understanding the shared role of expert and community-driven fact-checking in the global information ecosystem.
We find that at least 1 in 20 community notes rely explicitly on the work of professional fact-checkers, while this reliance is higher still for high-stakes topics such as health and politics. Our experiments also show that fact-checking is vital for debunking misleading content linked to broader narratives or conspiracy theories. These findings imply that high-quality community notes cannot be produced independently of professional fact-checking. They further suggest that the pressure on fact-checkers exerted by platforms and politicians by defunding and discrediting fact-checking organisations will have corrosive effects on the quality of notes and destructive implications for information integrity more widely.
2 Background
2.1 Community notes
Community moderation has been proposed as a means of addressing the scalability (Martel et al., 2024) and cross-partisanship trust (Flamini, 2019) challenges associated with fact-checking. Twitter/X’s Community Notes programme (piloted in 2021 and publicly launched in October 2022 (Twitter/X, 2021)) is a notable example of such a system. Any platform user may volunteer as a Community Notes contributor, although they must achieve a particular ‘rating impact score’ before they can write notes (Twitter/X, 2024b). Notes that achieve a ‘helpful’ rating appear underneath the post, explaining why the post is misleading (see Fig. 1). To be rated ‘helpful’, a note must receive similar levels of helpfulness rating from users with diverse viewpoints (Twitter/X, 2024a).
2.1.1 Characteristics of Community Notes
A small but growing body of work has analysed Twitter/X’s Community Notes dataset, focusing on the targets, sources, and limitations of notes.
Targets of notes. Community notes tend to focus on misleading posts from large accounts (Pilarski et al., 2024), focusing on posts that lack important content or present unverified claims as facts (Pröllochs, 2022; Drolsbach and Pröllochs, 2023).
Sources in notes. Analyses have indicated that notes were rated more helpful if they link to ‘trustworthy’ sources (Pröllochs, 2022) and that the majority of sources cited by notes were ‘trustworthy’ left-leaning news outlets. A recent study finds that 55% of URLs used in notes were related to news websites, 18% to research, 9% to social media, 9% to encyclopedic sources, but just 1.2% to fact-checking sources (Kangur et al., 2024).
Limitations of notes. Only 11% of submitted notes reach ‘helpful’ status (i.e., shown to users) by achieving a cross-perspective (Renault et al., 2024; Wirtschafter and Majumder, 2023), and the time frame for notes to reach the algorithm’s required agreement level (15.5 hours on average) limits its capacity to halt misinformation spread (Renault et al., 2024). Additional concerns about the notes’ efficacy highlight their indifference to the expertise needed for certain claims and reliance on subjective helpfulness rather than objective facts, free labour and inadequate support and guardrails regarding explicit content (Gilbert, 2025).
Our work provides novel insights into the targets, sources and limitations of community notes by shedding light on the relationship between notes and professional fact-checking. Namely, we study the extent to which fact-checking sources form the basis of note-writers’ efforts to counter misinformation and identify the strategies they employ.
2.1.2 Impact of Community Notes on misinformation spread
Posts identified by community notes as misleading have been found to attain less virality (reposts, quote tweets and replies) than non-misleading posts Drolsbach and Pröllochs (2023); Renault et al. (2024). Community notes have also been shown to increase the probability of tweet retractions and deletions and speed up the retraction process Gao et al. (2024); Renault et al. (2024). However, other studies have found less positive evidence; for example, that users’ followers, likes and engagement increase after their post receives a community note Wirtschafter and Majumder (2023). Curiously, one study claims that showing community notes on posts reduced the spread of misleading posts by an average of 61% Chuai et al. (2024a), while a more recent analysis by the same authors found no effect of community notes on engagement with misinformation Chuai et al. (2024c).
People shown community notes alongside misleading social media posts were more accurate in identifying misleading posts, and the notes were judged to be more trustworthy than context-free misinformation flags (e.g., "Checked by fact-checkers" or "Checked by other social media users"), regardless of (US-centric) political beliefs Drolsbach et al. (2024). People shown either community notes or related news article suggestions were both less likely to to believe and report misleading information compared to a control group: community notes were more effective in reducing belief and sharing intention for positive rumours, while articles were more effective for negative rumours Kankham and Hou (2024). On the other hand, displaying community notes leads users to post more negative and angry replies to misleading posts Chuai et al. (2024b), while crowd workers are also prone to cognitive biases, such as overestimating a statement’s truthfulness the more they liked its claimant, and general overconfidence in their ability to ascertain truthful statements Draws et al. (2022).
2.2 Professional fact-checking and community note practices
Although fact-checks and community notes share similarities in how they address misleading claims, they also differ in key elements of practice and techniques of persuasion and communication (Kankham and Hou, 2024). Fact-checking typically involves the analysis and verification of public claims (e.g., statements in news reports and social media). In addition to verifying claims, in recent years many fact-checking organisations have also assumed a wider role in combatting misinformation spread, conducting long-term investigative journalism projects and citizen media literacy programs (Juneja and Mitra, 2022). Professional fact-checkers in organisations signatory to the International Fact-Checking Network (IFCN) follow a rigorous set of principles and transparency commitments. [^2] https://www.ifcncodeofprinciples.poynter.org/the-commitments In contrast, any platform user can contribute to community notes under anonymity, and the rating approach relies on the ’wisdom of crowds’, with little oversight or transparency regarding the biases of the note-writers. Numerous studies have documented the structured workflow that fact-checkers follow: (i) claim selection; (ii) collecting evidence; (iii) deciding on a verdict; and (iv) writing the fact-checking article (Graves, 2017; Micallef et al., 2022; Warren et al., 2025). Fact-checking articles, which are subject to multiple rounds of editorial scrutiny, are more formal and standardised in tone and style than community notes, which vary considerably. Fact-checkers must rely on credible sources and evidence to convince the reader, while community note writers may employ a range of persuasion techniques, such as appeals to emotion or other logical fallacies. Moreover, community notes typically serve as direct rebuttals to misleading posts, while fact-checking articles may address a more general claim than is expressed in a specific post. Finally, fact-checking articles are a one-way exchange, while community notes represent a more horizontal and interactive dialogue between writer and recipient of the fact-check (Kankham and Hou, 2024).
Our work builds on current understanding of the relationship between professional fact-checking and amateur community moderation by examining the extent to which community note writers deploy the work of professional fact-checkers in their notes.
3 Dataset
Figure 2: The categories of links used by Community notes’ authors as a source.
We download files containing all community notes and their metadata from the official website, [^3] https://communitynotes.x.com/guide/en/under-the-hood/download-data which amounts to 1.5M notes authored between January 28th 2021 and January 6th 2025. Of these, a total of 135K are rated by the community as ‘Helpful’, 51K are rated ‘Not helpful’, and 1.3M are unpublished, i.e., did not receive enough community ratings to reach a verdict. See Fig. 6 in App. A for statistics.
We filter the notes as follows. First, we remove 526K non-English notes, which we identify by applying the language detection library fast-langdetect. [^4] https://github.com/LlmKira/fast-langdetect Then, we further filter 268K “unnecessary” notes—notes attached to tweets that are classified by the community as “not misleading”. Finally, to focus only on notes that are used to address misinformation, we filter out 44K notes that contain one of the words “ad”, “spam”, or “phishing”. Following these filtration steps, we are left with a dataset containing 664K notes.
The next step involves categorising the sources that the note authors use to support their claims. First, we use regex to extract all the URLs found in the notes. See Tab. 4 in App. A for a list of the top-100 most common domains. We classify each URL in our dataset of 664K notes into one of 13 categories (detailed in Fig. 2) using the pipeline described below.
  1. 1.
    Check whether the domain name of the URL is found in a manually curated list of domains of professional fact-checking organisations (See Tab. 3 in App. A for the full list). If so, classify the URL as “fact-checking”.
  2. 2.
    Otherwise, search for paraphrases of the word “fact-check” in the URL, [^5] These URLs mostly link to the fact-checking divisions of news outlets, e.g., https://apnews.com/article/fact-checking-909101991741 and classify it as “fact-checking” if a match was found.
  3. 3.
    Otherwise, check whether the domain name is found in Tab. 4, which the authors of this paper manually annotated.
  4. 4.
    Otherwise, use GPT-4666Version gpt-4o-2024-08-06. to classify the domain name into one of the 13 categories. LABEL:lst:prompt_link in App. B details the prompt we used.
  5. 5.
    Finally, if GPT-4 fails or outputs an unknown category, label the URL as “unknown”.
Using this pipeline, we successfully classify 95% of the URLs to one of the 13 categories.
Moreover, we further subsample the notes for performing the in-depth analysis required for answering RQ2 (§ 4.2). From the notes rated as ‘Helpful’, we sample 3.5K notes with a “Fact-checking” source and a random sample of 22K additional notes. We then used web crawling to scrape the text of the posts to which these notes were attached. We name this subset $\mathcal{S}_{\text{text}}$ for simplicity.
4 Analysis
Figure 3: Mean scores of community annotations of misleading posts.
We analyse the dataset prepared in § 3 to answer the two research questions defined in § 1.
4.1 RQ1: To what degree do community notes rely on fact-checkers?
According to Fig. 2, at least 5% of all English community notes contain an external link to professional fact-checkers. This number grows to 7% when only considering notes rated as ‘helpful’ (Fig. 7 in App. A). Conversely, only 1% of notes rated as ‘not helpful’ contain a fact-checking source (Fig. 8 in App. A). These figures are significantly larger than what was reported in previous studies (1.2% (Kangur et al., 2024)), possibly because Kangur et al. (2024) utilise a smaller dataset of fact-checking agencies and classify fact-checking divisions of popular journals as “news”. The results imply that notes incorporating fact-checking sources are generally considered more helpful.
We further assess whether notes with fact-checking sources are perceived to be of higher quality by analysing individual user ratings of notes both with and without such sources. Specifically, we collect user ratings for a balanced (i.e., including of a fact-checking source or not) sample of 20K notes rated by at least 50 users, and calculated the average ratings for the notes. As can be seen in Fig. 9 in App. A, community notes with fact-checking sources are generally rated higher than their counterparts. Interestingly, while notes with fact-checking links are more likely to be regarded as having a good source (higher HelpfulGoodSources), they are also more likely to be rated as notHelpfulSourcesMissingOrUnreliable. Tab. 5 in App. A contains a sample of such notes.
4.2 RQ2: What are the traits of posts and notes that rely on fact-checking sources?
Tweet Note
misleadingUnverifiedClaimAsFact
misleadingOutdatedInformation
misleadingFactualError
misleadingSatire
Fact Checking source
The NASA War Document is absolutely terrifying https://t.co/... misrepresenting a presentation by NASA scientist Dennis Bushnell, The lecture was not detailing plans by NASA to attack the world it was a lecture for defense industry professionals, and how defense tactics might rise to meet evolving threats in the future. https://leadstories.com/hoax-alert/2021/06/fact-check-the-future-is-now-is-not-a-nasa-war-document-plan-for-world-domination-and-phasing-out-of-humans.html
BREAKING NEWS: International Criminal Investigation calls on every public citizen to recommend indictments for Bill Gates, Anthony Fauci, Pfizer, BlackRock, Tedros and Christian Drosten for pushing everyone to receive the ineffective highly dangerous lethal experimental vaccines… Video has been fact-checked by USA Today, was found to be misleading, and promotes a conspiracy theory about COVID … https://ca.movies.yahoo.com/movies/fact-check-viral-video-promotes-204414488.html
1) California is RED. It is just because of the MASSIVE Election Fraud that stupid, brainwashed people believe Calif. is blue. Joe Biden won only in the SFO Bay area … The map shows the results of Reagan’s reelection in 1984, not Biden’s election in 2020. https://en.wikipedia.org/wiki/1984_United_States_presidential_election_in_California
Davis blows up \$100,000 fireworks in Kai Cenat setup During the Mr Beast Stream … The second photo is from a house fire in Atlanta in 2019. https://www.11alive.com/article/news/local/woodland-brook-drive-cause-of-house-fire/85-ecb7df9b-5f65-44e9-bf9d-8c162d36c334
@cnviolations I swear community notes are the only good thing Elon added since he bought Twitter. Community notes was first launched under former Twitter CEO Jack Dorsey in 2021 under the name of “Birdwatch”. The only thing Elon Musk did was that he renamed the feature to community notes. https://blog.twitter.com/en_us/topics/product/2021/introducing-birdwatch-a-community-based-approach-to-misinformation https://www.reuters.com/article/factcheck-elon-birdwatch-idUSL1N31Z2VG/
Thailand will become the first country to make the contract null and void, meaning that Pfizer will become responsible for all vaccine injuries … Thailand has no plans to void its Pfizer COVID vaccine contract, an official with the country’s National Vaccine Institute said. Thailand’s Department of Disease Control also rejected the claims as “fake news.” … https://apnews.com/article/fact-check-covid-vaccine-pfizer-thailand-203948163859
Hilarious tweets by footballers, A thread: 1. Virgil Van Dijk [Current Liverpool Captain] https://t.co/... Virgil Van Dijk did not tweet this, the tweet was made by a fan account in his name. https://www.pinkvilla.com/sports/fact-check-did-virgil-van-dijk-really-root-for-man-u-because-no-one-likes-liverpool-in-resurfaced-viral-tweet-1287250
Rob Reiner announces he’s on the Epstein Client List and Epstein Flight logs. What a fool! When a lawyer tells me to STFU, I STFU! https://t.co/... This is a digitally altered photo that might be misinterpreted even if used as a joke. The name Rob Reiner is misspelled, and the text is not on Reiner’s X timeline. https://twitter.com/robreiner?t=iqu43-NszIW5oOM_KqRSpw
Table 1: A sample of tweets, notes, and their community annotations, as well as whether the note contains a fact-checking link.
Figure 4: (a) strategies in debunking claims related to broader narratives. (b) the different ways in which fact-checking sources are used to debunk claims.
FC source
Conspi-racy 22% 11%
28% 39%
Table 2: Percentage of samples related to a broader narrative or conspiracy vs. have a fact-checking source.
We begin by performing a topic analysis, comparing topics of posts whose notes reference fact-checking sources to those citing other sources. To this end, we apply a strong zero-shot text classification model [^7] https://huggingface.co/r-f/ModernBERT-large-zeroshot-v1 with default settings. to our $\mathcal{S}_{\text{text}}$ subset by classifying spans of the form “Tweet:<POST TEXT>; Note <NOTE TEXT>” into one of 13 classes. The authors manually evaluated the quality of the classification results and considered it satisfactory. Notably (Fig. 5), fact-checking sources are more likely to be included in posts related to high-stakes issues such as health, science, and scams and less likely to be included in posts on tech or sports.
We then analyse annotations (binary attributes explaining the warrant for the note) by community note authors. Fig. 3 contains the full breakdown of annotations for notes with and without fact-checking sources. Notes containing a link to fact-checking sources are overrepresented in posts where unverified information is presented as a fact or when the post contains a factual error. Conversely, they are under-represented in posts with outdated information or satirical content. Tab. 1 contains a sample of such notes.
[…]
Figure 5: Distribution of notes’ topics, with and without a fact-checking source.
[…]
5 Conclusion
In this work, we annotate a large corpus of Twitter community notes with attributes such as topic, cited sources, and whether they refute claims tied to broader misinformation narratives. We find that effective community moderation depends on professional fact-checking to an extent far greater than previously reported. We find that community notes linked to broader narratives or conspiracy theories are particularly reliant on fact-checking.
Our results reveal that community notes and professional fact-checking are deeply interconnected—fact-checkers conduct in-depth research beyond the reach of amateur platform users, while community notes publicise their work. The move by platforms to end their partnerships and funding for fact-checking organisations will hinder their ability to fact-check and pursue investigative journalism, which community note writers rely on. This, in turn, will limit the efficacy of community notes, especially for high-stakes claims tied to broader narratives or conspiracies.
Limitations
The main limitations of our work concern the characteristics of the dataset we analyse. First, we restrict our analysis to notes written in English, excluding over half a million notes in other languages. This decision was made to avoid potential noise and biases arising from the authors’ unfamiliarity with public discourse in different regions and reliance on machine translation. In future work, we aim to extend our analysis to other languages.
Moreover, except for a small subset of notes, we did not have access to the original tweets they were written for. Even when the tweet text was available, many contained non-text media, were written in internet vernacular that was challenging to interpret, or lacked important context. These factors limit the accuracy and effectiveness of our models and analysis.
Finally, due to resource constraints, our manual annotation study was limited to a relatively small sample of tweets and notes. In future work, we wish to utilise crowd workers to not only annotate a larger dataset but also increase the diversity and perspective of the annotators.
Broader Impact and Ethical Considerations
Given that this work analyses real-world posts, ethical concerns may arise from using this data for research purposes. Posts from non-protected accounts and Community Notes on Twitter/X are publicly available, however, we acknowledge that they may contain sensitive personal information. To minimise any breach of anonymity and privacy, we anonymised links to individual accounts, and we do not publicly release this information. We do not analyse the posts or notes by individual users, and instead examine aggregated data in the form of topics and sources cited.
Although the Community Notes dataset represents attempts to curb harmful misinformation and conspiracies, given the intense partisanship involved (Allen et al., 2022; Draws et al., 2022), as well as the explicit content of some claims, some instances may be considered offensive. We also acknowledge that our own perspectives and biases as authors shape the impact of our findings in certain ways. For example, as mentioned in the previous section, we were unable to analyse non-English posts in-depth, so our conclusions are likely somewhat focused on discourse in the Anglosphere (e.g., the US, UK, Ireland, Canada, Australia, New Zealand etc.). Furthermore, although we based our criteria for conspiracy theories on well-established sources, e.g., AP News, FactCheck.org, the European Commission, and identified conspiratorial narratives from both left- and right-wing sources, our own perspectives (i.e., as scientists from Western countries) may also have impacted what we considered to be conspiracy theories.
Acknowledgements
[Uncaptioned image][Uncaptioned $\end{array}$ This research was co-funded by the European Union (ERC, ExplainYourself, 101077481), by the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101135671 (TrustLLM), and by the Pioneer Centre for AI, DNRF grant number P1.
Appendix A Additional Material
Figure 6: A histogram of the number of community notes written every month and their rating (helpful, not helpful, or needs more data. The grey vertical line (December 2022) indicates the date when the community notes became visible worldwide.
Figure 7: The categories of links used by Community notes’ authors as a source, filtering for notes rated as “helpful”.
Figure 8: The categories of links used by Community notes’ authors as a source, filtering for notes rated as “not helpful”.
Figure 9: Community ratings of notes with and without fact-checking source.
Name URL Language Region/domain
Lead stories leadstories.com English Global
AFP Factuel factuel.afp.com French Global
AAP FactCheck aap.com.au/factcheck English Australia
Full Fact fullfact.org English Global
Science Feedback science.feedback.org English Science
Politifact politifact.com English, Spanish USA
HoaxEye hoaxeye.wordpress.com English Images
Logically Facts logicallyfacts.com Multiple Europe/India
FactCheckNI factcheckni.org English North Ireland
DFRLab dfrlab.org English Global
FactReview factreview.gr Greek Global
Lupa lupa.uol.com.br/jornalismo Portuguese Global
Check your fact checkyourfact.com English Global
Climate feedback climatefeedback.org English Climate
Factcheck factcheck.org English USA
Health feedback healthfeedback.org English Health
Snopes snopes.com English US
aosfatos aosfatos.org Portuguese Global
Demagog demagog.org.pl/fake_news Polish Poland
FakeReporter fakereporter.net Hebrew Israel
litmus factcheck litmus-factcheck.jp Japanese Japan
Climate Feedback climatefeedback.org English Global
AFP factcheck.afp.com English Global
USA Today usatoday.com/story/news/factcheck English USA
Statesman statesman.com English USA
Dallas News dallasnews.com/news/politifact English USA
Google Fact Check toolbox.google.com/factcheck English Global
MediaBias/FactCheck mediabiasfactcheck.com English Global
MedDMO meddmo.eu English, Greek Greece, Cyprus, Malta
Poynter poynter.org/fact-checking English USA
Newsmeter newsmeter.in/fact-check English, Tamil India
Africa Check africacheck.org English Africa
Fact Crescendo India english.factcrescendo.com English India
Factseeker factseeker.lk English Sri Lanka
Fact Crescendo Thailand thailand.factcrescendo.com Thai Thailand
Fact Crescendo Afghanistan afghanistan.factcrescendo.com Persian Afghanistan
Only Fact onlyfact.in English India
Factly factly.in English India
Fact Crescendo Sri Lanka srilanka.factcrescendo.com Sinhala Sri Lanka
Fact Crescendo Cambodia cambodia.factcrescendo.com Cambodian Cambodia
Becid becid.eu Baltic langs Baltic
Fact Hunt facthunt.in English India
Tec Arp techarp.com English Global (based in Malaysia)
10 news 10news.com/news/fact-or-fiction English USA
RMIT Fact Check rmit.edu.au English Australia
Gigafact gigafact.org English USA
Ayupp ayupp.com/fact-check English India
The Journal thejournal.ie English Ireland
Table 3: List of professional fact-checking organisations and their URLs.
Domain Category Domain Category
x.com social media thehill.com news
twitter.com social media amp.theguardian.com news
youtube.com social media whitehouse.gov government
youtu.be social media news.sky.com news
un.org organisation merriam-webster.com reference
u.today news techarp.com news
t.co social media cbc.ca news
snopes.com fact checking politifact.com fact checking
en.m.wikipedia.org reference pbs.org commercial
en.wikipedia.org reference telegraph.co.uk news
google.com search engine businessinsider.com news
instagram.com social media time.com news
britannica.com reference justice.gov government
reuters.com news cnbc.com news
bbc.co.uk news wsj.com news
apnews.com news sciencedirect.com academic
bbc.com news msn.com news
nytimes.com news statista.com reference
theguardian.com news business.x.com commercial
vice.com news amp.cnn.com news
usatoday.com news congress.gov government
factcheck.org fact checking factcheck.afp.com fact checking
cnn.com news yahoo.com search engine
washingtonpost.com news timesofindia.indiatimes.com news
ncbi.nlm.nih.gov academic thelancet.com academic
nbcnews.com news hrw.org organisation
help.twitter.com reference healthfeedback.org fact checking
cdc.gov government fda.gov government
npr.org news m.youtube.com social media
forbes.com news law.cornell.edu academic
newsweek.com news medium.com blog post
fullfact.org fact checking healthfeedback.org fact checking
dailymail.co.uk news who.int organisation
cbsnews.com news haaretz.com news
web3antivirus.io database axios.com news
timesofisrael.com news mayoclinic.org commercial
help.x.com reference nejm.org academic
nypost.com news scienceexchange.caltech.edu academic
aljazeera.com news indiatoday.in news
reddit.com social media bloomberg.com news
independent.co.uk news pewresearch.org academic
usgs.gov academic jamanetwork.com academic
abcnews.go.com news leadstories.com news
nature.com academic dictionary.cambridge.org reference
gov.uk government jpost.com news
web.archive.org database archive.ph database
foxnews.com news healthline.com commercial
tiktok.com social media abc.net.au news
edition.cnn.com news france24.com news
Table 4: List of top 100 most common domains found in the community notes dataset, and their categorization.
ID summary
0 This claim ruled mostly false. https://www.politifact.com/factchecks/2020/may/07/facebook-posts/facebook-post-cites-doctors-widely-disputed-calcul/
1 The RedState article claims “the shots do not stop transmission of the virus. This is false. ”“Vaccines provide significant protection from ’getting it’ – infection – and ’spreading it’ – transmission – even against the delta variant.” Source: https://www.usatoday.com/story/news/factcheck/2021/11/17/fact-check-covid-19-vaccines-protect-against-infection-transmission/6403678001/
2 There is no proof of this, the photo is real, it’s not the last photo of the child. But snoops say there is a tenuous link the parents used the same law firm to represent them as Maxwell https://www.snopes.com/fact-check/ghislaine-maxwell-jonbenet-ramsey/
3 unfounded https://www.snopes.com/fact-check/ashley-biden-diary-afraid/
4 The mRNA vaccine does not cause cancer: https://www.factcheck.org/2024/05/still-no-evidence-covid-19-vaccination-increases-cancer-risk-despite-posts/
5 Many of the details in this popular essay are inaccurate and too numerous to list here. The essay was fact checked by Snopes in 2005: https://www.snopes.com/fact-check/the-price-they-paid/
6 POLITIFACT - rates False. The report analysed a small sample of 128 temp stations out of several thousand volunteer-run stations, then extrapolated results. NOAA uses 2 programs to record daily temps. The report did not look at the 900 more sophisticated automated stations. https://www.politifact.com/factchecks/2022/aug/19/facebook-posts/fact-checking-talking-point-about-corrupted-climat/
7 There is no verifiable evidence of campaign espionage in either the 2020 or the 2016 presidential elections. https://www.snopes.com/fact-check/obama-spying-trump-campaign/ https://www.washingtonpost.com/politics/2019/05/06/whats-evidence-spying-trumps-campaign-heres-your-guide/
8 Ladapo did get caught altering COVID vaccine study findings. Ladapo replaced the language from an earlier study draft that found no significant risk from COVID vaccines, to then state there was a high risk https://healthfeedback.org/claimreview/analysis-florida-department-health-surgeon-general-joseph-ladopo-contains-multiple-methodological-problems-covid-19-mrna-vaccines/ https://healthexec.com/topics/clinical/COVID-19/florida-surgeon-general-altered-covid-19-study-findings
Table 5: Examples of community notes containing fact-checking sources that are rated as having notHelpfulSourcesMissingOrUnreliable.
This section details additional results or material referenced from the paper’s main body.
Fig. 6 A histogram of the number of community notes written every month and their rating (helpful, not helpful, or needs more data).
Fig. 7 The categories of links used by Community notes’ authors as a source, filtering for notes rated as “helpful”.
Fig. 8 The categories of links used by Community notes’ authors as a source, filtering for notes rated as “not helpful”.
Fig. 9 Community ratings of notes with and without fact-checking source.
Tab. 3 List of professional fact-checking organisations and their URLs.
Tab. 3 List of top 100 most common domains found in the community notes dataset, and their categorization.
Tab. 5 Examples of community notes containing fact-checking sources that are rated as having notHelpfulSourcesMissingOrUnreliable.
Appendix B Reproducibility
Listing 1: The prompt used to classify URLs into categories.
SYSTEM PROMPT
You are a professional IT system who has a vast knowledge of the internet and its content. Your goal is simple, but very important: Classify URLs into categories. Choose only from the provided categories!
USER PROMPT
Read the following URLs.
Your goal is to categorize each url into one of the pre-defined categories.
Chose from the following list of categories:
Categories =
"social media", # Social media sites like Facebook, Twitter, Youtube etc.
"news", # Websites of news outlets or other organisations that report current events, such as the nytimes, the guardian, etc.
"government", # Government agencies and organisations, as well as websites related to policies and guidelines, such as the CDC, department of education, FDA, etc.
"academic", # Academic sources, journals, and magazines, such as pubmed, nature, sciencedirect, etc.
"blog post", # Independent blog posts about various topics, including cooking, travel, home improvement, fandom, reviews, etc.
"fact checking", # professional fact checking organisations
"database", # Public databases such as google drive, archive.com, dropbox, etc.
"commercial", # Webpages of commercial organisations such as BMW, Delta, Nike, etc.
"reference", # Public resources such as encyclopedias, dictionaries, advocacy sources, guides, DIYs, statistics, religious sources, travel information, usage guidelines, Q&As, terms of services, etc.
"organisation", # non-commercial and non-government organisations such as WHO, the UN, Greenpeace, LA-Lakers, etc.
"other", # Any other website that does not fit into one of the previous categories.
"unknown", # if it is impossible to determine the category of the webpage.
Output format example:
id: <ID>,
url: <URL>,
category: <CATEGORY>,
URLs:
<URLS>
Listing 2: The prompt used to classify tweets and notes into broader narratives and conspiracy theories.
SYSTEM PROMPT
You are a professional fact-checker who specializes in analyzing misinformation spread on social media.
Your goal is to analyse a tweet and a community note written about the tweet and decide whether the tweet spread misinformation related to a known conspiracy theory or a misleading wider narrative, and if so, which one is it.
USER PROMPT
[…]
Include your reasoning. Output the results as a json file. If a tweet does not relate to a conspiracy theory or a misleading wider narrative, output "none" in the json.
- Tweets *do not* discuss a wider narrative if the misleading information is tied to a specific singular event that is not connected to major topics on the public discourse.
They do discuss a wider narrative if the misleading information is tied to a known conspiracy theory or to major topics on the public discorse.
Chose from the following list of theories and wider narrative:
CONSPIRACY_THEORIES =
September 11,
October 7,
the great replacement,
COVID was intentionally spread,
the COVID outbreak is fake,
2020 election fraud,
vaccines cause autism,
5G towers,
Russian invasion of Ukraine,
flat earth,
chemtrails,
Q-Anon and deep state,
Epstein files,
Barack Obama was not born in the USA,
Michelle Obama is a man,
LGBT grooming,
fluorite in the water,
climate change,
Holocaust denial,
Hunter Biden and Ukraine,
other,
Output format example:
id: <ID>,
is_related_to_conspiracy: <True/False>,
conspiracy: <CONSOIRACY (or None)>,
reasoning: <REASONING>\
Tweets and notes:
<TWEETS_AND_NOTED>
LABEL:lst:prompt_link The prompt used to classify URLs into categories.
LABEL:lst:prompt_conspiracy The prompt used to classify tweets and notes into broader narratives and conspiracy theories.
B.1 Manual Annotation Setup
Figure 10: Our annotation setup.
[…]
Broader narrative
[…]
Discredit source of claim
If the community note describes the source shared by the original post as non-credible.
Add missing context
If the community note provides some missing context to refute a claim.
Highlight AI generated
If the community note claims that the post shared AI-generated content.
Highlight edited media
If the community note claims that the post shared some media that was edited (edited with Photoshop, the clip was cut, etc.).
Link to direct source
If the community note shares a link to a source where an entity says that a claim made about them is false.
Link official source
If the community note shares a link to an official source such as a government website.
Link scientific source
If the community note shares a link to some scientific article or website.
Link world knowledge
If the community note shares a link to some reference resources such as Wikipedia.
Link fact-checking
If the community note shares a link to a professional fact-checking organisation.
In-note fact-checking
If the community note performs an in-note fact-check by cross-referencing several sources and constructing a compelling argument.